TEX Mathematical Symbols

Paul Taylor 1 June 1993

The following mathematical symbols are available in plain TEX, ^{†}mssymb.tex (or amssymb.sty) and ^{♯}stmaryrd.sty. Multiple column output courtesy of Frank Mittelbach’s multicol.sty.

\aleph ℵ
\alpha α
\amalg ⨿
\angle ∠
\approx ≈
\approxeq^{†} ≊

\arccos arccos
\arcsin arcsin
\arctan arctan
\arg arg

\Arrownot^{♯}
\arrownot^{♯}
\ast ∗
\asymp ≍
\backepsilon^{†} ϶
\backprime^{†} ‵
\backsim^{†} ∽
\backsimeq^{†} ⋍
\backslash \
\baro^{♯} ⏀
\barwedge^{†} ⊼
\bbslash^{♯} ⑊
\because^{†} ∵
\beta β
\beth^{†} ℶ
\between^{†} ≬
\bigbox^{♯} □
\bigcap ⋂
\bigcirc ◯
\bigcup ⋃
\bigcurlyvee^{♯} ⋎
\bigcurlywedge^{♯} ⋏
\biginterleave^{♯} ⫼
\bignplus^{♯}
\bigodot ⨀
\bigoplus ⨁
\bigotimes ⨂
\bigparallel^{♯} ∥
\bigsqcap^{♯} ⨅
\bigsqcup ⨆

\bigstar^{†} ⋆
\bigtriangledown ▽
\bigtriangledown^{♯} ▽
\bigtriangleup △
\bigtriangleup^{♯} △
\biguplus ⨄
\bigvee ⋁
\bigwedge ⋀
\binampersand^{♯} &
\bindnasrepma^{♯} ⅋
\blacklozenge^{†} ♦
\blacksquare^{†} ■
\blacktriangle^{†} ▲
\blacktriangledown^{†} ▼
\blacktriangleleft^{†} ◀
\blacktriangleright^{†} ▶
  \bmod mod
   \bot ⊥
\bowtie ▷◁
\boxast^{♯} ⧆
\boxbar^{♯} ◫
\boxbox^{♯} ⧈
\boxbslash^{♯} ⧅
\boxcircle^{♯} ⧇
\boxdot^{♯} ⊡
\boxdot^{†} ⊡
\boxempty^{♯} □
\boxminus^{†} ⊟
\boxplus^{†} ⊞
\boxslash^{♯} ⧄
\boxtimes^{†} ⊠
\bullet •
\Bumpeq^{†} ≎
\bumpeq^{†} ≏
   \Cap^{†} ⋒
   \cap ∩
  \cdot ·
 \cdotp ·
 \cdots · · ·
\centerdot^{†} ▪

                         \checkmark^{†} ✓
                               \chi χ
                              \circ ∘
                            \circeq^{†} ≗
                   \circlearrowleft^{†} ⟲
                  \circlearrowright^{†} ⟳
                        \circledast^{†} ⊛
                       \circledcirc^{†} ⊚
                       \circleddash^{†} ⊝
                          \circledR^{†} ®
                          \circledS^{†} Ⓢ
                          \clubsuit ♣
                             \colon :
                        \complement^{†} ∁
                              \cong   ≅
                            \coprod ∐
                         \copyright   \begin{matrix}c \\ ◯ \end{matrix}

\cos cos
\cosh cosh
\cot cot
\coth coth
\csc csc

                               \Cup^{†} ⋓
                               \cup ∪
                       \curlyeqprec^{†} ⋞
                       \curlyeqsucc^{†} ⋟
                          \curlyvee^{†} ⋎
                 \curlyveedownarrow^{♯}
                   \curlyveeuparrow^{♯}
                        \curlywedge^{†} ⋏
               \curlywedgedownarrow^{♯}
                 \curlywedgeuparrow^{♯}
                    \curvearrowleft^{†} ↶
                   \curvearrowright^{†} ↷
                               \dag †
                            \dagger †
                            \daleth^{†} ℸ
                             \dashv ⊣
                              \ddag ‡

\ddagger ‡
\ddots ...

\deg deg

\Delta Δ
\delta δ

\det det

\diamond ⋄
\diamondsuit ♢
\digamma^{†} 𝟋

\dim dim

  \div ÷
\divideontimes^{†} ⋇
\Doteq^{†} ≑
\doteq   ≐
\doteqdot^{†} ≑
\dotplus^{†} ∔
 \dots . . .
\doublebarwedge^{†} ⩞
\doublecap^{†} ⋒
\doublecup^{†} ⋓
\Downarrow ⇓
\downarrow ↓
\downdownarrows^{†} ⇊
\downharpoonleft^{†} ⇃
\downharpoonright^{†} ⇂
  \ell ℓ
\emptyset ∅
\epsilon ϵ
\eqcirc^{†} ≖

\eqsim^{†}

\eqslantgtr^{†} ⪖
\eqslantless^{†} ⪕
\equiv ≡
  \eta η

  \eth^{†}

\exists ∃

\exp exp

\fallingdotseq^{†} ≒
\fatbslash^{♯}
\fatsemi^{♯} ⨟
\fatslash^{♯}
 \flat ♭
\forall ∀
\frown ⌢
\Gamma Γ
\gamma γ

\gcd gcd

  \geq ≥
 \geqq^{†} ≧
\geqslant^{†} ⩾
   \gg ≫
  \ggg^{†} ≫
\gggtr^{†} ≫

                            \gimel^{†} ℷ
                         \gnapprox^{†} ⪊
                             \gneq^{†} ⪈
                            \gneqq^{†} ≩
                            \gnsim^{†} ⋧
                        \gtrapprox^{†} ⪆
                           \gtrdot^{†} ⋗
                        \gtreqless^{†} ⋛
                       \gtreqqless^{†} ⪌
                          \gtrless^{†} ≷
                           \gtrsim^{†} ≳
                        \gvertneqq^{†} ≩
                             \hbar ℏ
                             \hbar^{†} ℏ
                        \heartsuit ♡

\hom hom

                           \hslash^{†} ℏ
                              \iff ⟺
                               \Im ℑ
                            \imath ı
                               \in ∈

\inf inf

                            \infty ∞
                           \inplus^{♯}
                         \intercal^{†} ⊺
                       \interleave^{♯} ⫴
                            \intop ∫
                             \iota ι
                            \jmath ȷ
                            \kappa κ

\ker ker

                           \Lambda Λ
                           \lambda λ
                           \langle ⟨
                             \Lbag^{♯} ⟅
                             \lbag^{♯} ⟅
                           \lbrace {
                            \lceil ⌈
                            \ldotp .
                            \ldots . . .
                        \Leftarrow ⇐
                        \leftarrow ←
                    \leftarrowtail^{†} ↢
                \leftarrowtriangle^{♯} ⇽
                  \leftharpoondown ↽
                    \leftharpoonup ↼
                   \leftleftarrows^{†} ⇇
                   \Leftrightarrow ⇔
                   \leftrightarrow ↔
                 \leftrightarroweq^{♯}
                  \leftrightarrows^{†} ⇆
           \leftrightarrowtriangle^{♯} ⇿
                \leftrightharpoons^{†} ⇋

                                         \leftrightsquigarrow^{†} ↭
                                                   \leftslice^{♯} ⪦
                                              \leftthreetimes^{†} ⋋
                                                         \leq ≤
                                                        \leqq^{†} ≦
                                                    \leqslant^{†} ⩽
                                                  \lessapprox^{†} ⪅
                                                     \lessdot^{†} ⋖
                                                   \lesseqgtr^{†} ⋚
                                                  \lesseqqgtr^{†} ⪋
                                                     \lessgtr^{†} ≶
                                                     \lesssim^{†} ≲
                                                      \lfloor ⌊

\lg lg

                                                       \lhook
                                                   \lightning^{♯} ⚡

\lim lim
\liminf lim inf
\limsup lim sup

                                                          \ll ≪
                                                   \llbracket^{♯} ⟦
                                                      \llceil^{♯}
                                                    \llcorner^{†} ⌞
                                                  \Lleftarrow^{†} ⇚
                                                     \llfloor^{♯}
                                                         \lll^{†} ≪
                                                      \llless^{†} ≪
                                               \llparenthesis^{♯} ⦇

\ln ln

                                                    \lnapprox^{†} ⪉
                                                        \lneq^{†} ⪇
                                                       \lneqq^{†} ≨
                                                       \lnsim^{†} ⋦

\log log

                                                \Longarrownot^{♯}
                                                \longarrownot^{♯}
                                               \Longleftarrow ⟸
                                               \longleftarrow ⟵
                                          \Longleftrightarrow ⟺
                                          \longleftrightarrow ⟷
                                                \Longmapsfrom^{♯} ⟸
                                                \longmapsfrom^{♯} ⟵
                                                  \Longmapsto^{♯} ⟹
                                                  \longmapsto ⟼
                                              \Longrightarrow ⟹
                                              \longrightarrow ⟶
                                               \looparrowleft^{†} ↫
                                              \looparrowright^{†} ↬
                                                     \lozenge^{†} ♢
                                                    \lrcorner^{†} ⌟
                                                         \Lsh^{†} ↰
                                                      \ltimes^{†} ⋉
                                                   \lvertneqq^{†} ≨

\maltese^{†} ✠
\Mapsfrom^{♯} ⤆
\mapsfrom^{♯} ↤
\Mapsfromchar^{♯}
\mapsfromchar^{♯}
\Mapsto^{♯} ⤇
\Mapstochar^{♯}
\mapstochar

\max max

\measuredangle^{†} ∡
\merge^{♯} ⩕
\mho^{†} ℧
\mid |

\min min

\minuso^{♯} ⦵
\models ⊧
\moo^{♯}
\mp ∓
\mu μ
\multimap^{†} ⊸
\nabla ∇
\napprox ≇
\natural ♮

\ncong^{†}

\nearrow ↗
\neg ¬
\neq ≠
\nexists^{†} ∄
\ngeq^{†} ≱
\ngeqq^{†} ≧̸
\ngeqslant^{†} ⩾̸
\ngtr^{†} ≯
\ni ∋
\niplus^{♯}
\nLeftarrow^{†} ⇍
\nleftarrow^{†} ↚
\nLeftrightarrow^{†} ⇎
\nleftrightarrow^{†} ↮
\nleq^{†} ≰
\nleqq^{†} ≦̸
\nleqslant^{†} ⩽̸
\nless^{†} ≮
\nmid^{†} ∤
\nnearrow^{♯} ↗
\nnwarrow^{♯} ↖
\not
\notin   ∉
\nparallel^{†} ∦
\nplus^{♯}
\nprec^{†} ⊀
\npreceq^{†} ⪯̸
\nRightarrow^{†} ⇏
\nrightarrow^{†} ↛

\nshortmid^{†} ∤
\nshortparallel^{†} ∦
  \nsim^{†} ≁
\nsubseteq^{†} ⊈
\nsubseteqq^{†} ⫅̸
 \nsucc^{†} ⊁
\nsucceq^{†} ⪰̸
\nsupseteq^{†} ⊉
\nsupseteqq^{†} ⫆̸
\ntriangleleft^{†} ⋪
\ntrianglelefteq^{†} ⋬
\ntrianglelefteqslant^{♯} ⋬
\ntriangleright^{†} ⋫
\ntrianglerighteq^{†} ⋭
\ntrianglerighteqslant^{♯} ⋭
    \nu ν
\nVDash^{†} ⊯
\nVdash^{†} ⊮
\nvDash^{†} ⊭
\nvdash^{†} ⊬
\nwarrow ↖
  \obar^{♯} ⌽
\oblong^{♯} □
\obslash^{♯} ⦸
  \odot ⊙
\ogreaterthan^{♯} ⧁
\ointop ∮
\olessthan^{♯} ⧀
 \Omega Ω
 \omega ω
\ominus ⊖
 \oplus ⊕
\oslash ⊘
\otimes ⊗
  \ovee^{♯}
\owedge^{♯}

\P ¶

\parallel ∥
\partial ∂
  \perp ⊥
   \Phi Φ
   \phi ϕ
    \Pi Π
    \pi π
\pitchfork^{†} ⋔
    \pm ±

\Pr Pr

  \prec ≺
\precapprox^{†} ≾
\preccurlyeq^{†} ≼
\preceq ⪯
\precnapprox^{†} ⪹
\precneqq^{†} ⪵
\precnsim^{†} ⋨

                           \precsim^{†} ≾
                             \prime ′
                              \prod ∏
                            \propto ∝
                               \Psi Ψ
                               \psi ψ
                            \rangle ⟩
                              \Rbag^{♯} ⟆
                              \rbag^{♯} ⟆
                            \rbrace }
                             \rceil ⌉
                                \Re ℜ
                            \Relbar =
                            \relbar −
                       \restriction^{†} ↾
                            \rfloor ⌋
                               \rho ρ
                             \rhook
                        \Rightarrow ⇒
                        \rightarrow →
                    \rightarrowtail^{†} ↣
                \rightarrowtriangle^{♯} ⇾
                  \rightharpoondown ⇁
                    \rightharpoonup ⇀
                   \rightleftarrows^{†} ⇄
                 \rightleftharpoons ⇌
                 \rightleftharpoons^{†} ⇌
                  \rightrightarrows^{†} ⇉
                        \rightslice^{♯} ⪧
                   \rightsquigarrow^{†} ⇝
                   \rightthreetimes^{†} ⋌
                      \risingdotseq^{†} ≓
                        \rmoustache Γ
                         \rrbracket^{♯} ⟧
                            \rrceil^{♯}
                           \rrfloor^{♯}
                       \Rrightarrow^{†} ⇛
                     \rrparenthesis^{♯} ⦈
                               \Rsh^{†} ↱
                            \rtimes^{†} ⋊

\S §

                           \searrow ↘

\sec sec

                          \setminus \
                             \sharp ♯
                    \shortdownarrow^{♯} ↓
                    \shortleftarrow^{♯} ←
                          \shortmid^{†} ∣
                     \shortparallel^{†} ∥
                   \shortrightarrow^{♯} →
                      \shortuparrow^{♯} ↑
                             \Sigma Σ
                             \sigma σ
                               \sim ∼

   \simeq ≃

\sin sin
\sinh sinh

\smallfrown^{†} ⌢
\smallint ∫
\smallsetminus^{†} ∖
\smallsmile^{†} ⌣
   \smile ⌣
\spadesuit ♠
\sphericalangle^{†} ∢
   \sqcap ⊓
   \sqcup ⊔
    \sqrt √
\sqsubset^{†} ⊏
\sqsubseteq ⊑
\sqsupset^{†} ⊐
\sqsupseteq ⊒
  \square^{†} □
\ssearrow^{♯} ↘
  \sslash^{♯} ⫽
\sswarrow^{♯} ↙
    \star ⋆
  \Subset^{†} ⋐
  \subset ⊂
\subseteq ⊆
\subseteqq^{†} ⫅
\subsetneq^{†} ⊊
\subsetneqq^{†} ⫋
\subsetplus^{♯} ⪿
\subsetpluseq^{♯}
    \succ ≻
\succapprox^{†} ≿
\succcurlyeq^{†} ≽
  \succeq ⪰
\succnapprox^{†} ⪺
\succneqq^{†} ⪶
\succnsim^{†} ⋩
 \succsim^{†} ≿
     \sum ∑

\sup sup

  \Supset^{†} ⋑
  \supset ⊃
\supseteq ⊇
\supseteqq^{†} ⫆
\supsetneq^{†} ⊋
\supsetneqq^{†} ⫌

                          \supsetplus^{♯} ⫀
                        \supsetpluseq^{♯}
                                \surd √
                             \swarrow ↙
                          \talloblong^{♯} ⫾

\tan tan
\tanh tanh

                                 \tau τ

\TeX TEX

                           \therefore^{†} ∴
                               \Theta Θ
                               \theta θ
                         \thickapprox^{†} ≈
                            \thicksim^{†} ∼
                               \times ×
                                 \top ⊤
                            \triangle △
                        \triangledown^{†} ▽
                        \triangleleft ◁
                      \trianglelefteq^{†} ⊴
                 \trianglelefteqslant^{♯} ⊴
                           \triangleq^{†} ≜
                       \triangleright ▷
                     \trianglerighteq^{†} ⊵
                \trianglerighteqslant^{♯} ⊵
                    \twoheadleftarrow^{†} ↞
                   \twoheadrightarrow^{†} ↠
                            \ulcorner^{†} ⌜
                             \Uparrow ⇑
                             \uparrow ↑
                         \Updownarrow ⇕
                         \updownarrow ↕
                       \upharpoonleft^{†} ↿
                      \upharpoonright^{†} ↾
                               \uplus ⊎
                             \Upsilon Υ
                             \upsilon υ
                          \upuparrows^{†} ⇈
                            \urcorner^{†} ⌝
                          \varbigcirc^{♯} ◯
                        \varcopyright^{♯} \begin{matrix}c \\ ◯ \end{matrix}
                         \varcurlyvee^{♯} ⋎
                       \varcurlywedge^{♯} ⋏
                          \varepsilon ε
                            \varkappa^{†} κ
                          \varnothing^{†} ∅
                             \varoast^{♯} ⊛

                                                        \varobar^{♯} ⦶
                                                     \varobslash^{♯} ⍉
                                                     \varocircle^{♯} ⊚
                                                        \varodot^{♯} ⊙
                                                \varogreaterthan^{♯} ⧁
                                                   \varolessthan^{♯} ⧀
                                                      \varominus^{♯} ⊖
                                                       \varoplus^{♯} ⊕
                                                      \varoslash^{♯} ⊘
                                                      \varotimes^{♯} ⊗
                                                        \varovee^{♯}
                                                      \varowedge^{♯}
                                                         \varphi φ
                                                          \varpi ϖ
                                                      \varpropto^{†} ∝
                                                         \varrho ϱ
                                                       \varsigma ς
                                                   \varsubsetneq^{†} ⊊
                                                  \varsubsetneqq^{†} ⫋
                                                   \varsupsetneq^{†} ⊋
                                                  \varsupsetneqq^{†} ⫌
                                                       \vartheta ϑ
                                                       \vartimes^{♯} ⨉
                                                    \vartriangle^{†} △
                                                \vartriangleleft^{†} ◁
                                               \vartriangleright^{†} ▷
                                                          \Vdash^{†} ⊩
                                                          \vDash^{†} ⊨
                                                          \vdash ⊢
                                                          \vdots   ⋮
                                                            \vee ∨
                                                         \veebar^{†} ⊻
                                                           \Vert ∥
                                                           \vert |
                                                         \Vvdash^{†} ⊪
                                                          \wedge ∧
                                                             \wp ℘
                                                             \wr ≀
                                                             \Xi Ξ
                                                             \xi ξ
                                                          \Ydown^{♯}
                                                            \yen^{†} ¥
                                                          \Yleft^{♯}
                                                         \Yright^{♯}
                                                            \Yup^{♯} ⅄
                                                           \zeta ζ